Two blocks of mass 0.1 kg and 0.2 kg approach each other on a horizontal plane at velocities of 0.4 and 1 m/s respectively. If the blocks collide and remain together, calculate the joint velocity after collision.


1
Expert's answer
2021-05-07T06:40:23-0400

The total momentum before the collision and after the collision is conserved

"p_1+p_2 = p_3"

"m_1v_1+m_2v_2=m_3v_3"

"\displaystyle v_3 = \frac{m_1v_1+m_2v_2}{m_3} = \frac{0.1 \cdot 0.4 - 0.2 \cdot 1}{0.1+0.2} = -0.533" m/s.

The minus sign in the formula above appears because the velocity of the blocks had different directions. The minus sign of "v_3" just means that the joint block will move in the direction of the second block.


Answer: 0.533 m/s.